# Analysis of reactor burnup simulation uncertainties for antineutrino spectrum prediction

### v2

@@ Abstract
Nuclear reactors are a source of electron antineutrinos due to the presence of unstable fission products that undergo $\beta^{-}$ decay. They will be exploited by the JUNO experiment to determine the neutrino mass ordering and to get very precise measurements of the neutrino oscillation parameters. This requires the reactor antineutrino spectrum to be characterized as precisely as possible both through high resolution measurements, as foreseen by the TAO experiment, and detailed simulation models. In this paper we present a benchmark analysis utilizing Serpent Monte Carlo simulations in comparison with real pressurized water reactor spent fuel data. Our objective is to study the accuracy of fission fraction predictions as a function of different reactor simulation approximations. Then, using the BetaShape software, we construct fissile antineutrino spectra using the summation method, thereby assessing the influence of simulation uncertainties on the reactor antineutrino spectrum.
@@ 1 Introduction
which has a $\bar{\nu}_{e}$ energy threshold at 1.8 MeV.
One of the main goals of the JUNO experiment is to determine the neutrino mass ordering by analyzing the pattern of $\bar{\nu}_{e}$ oscillations at medium baseline. This experimental measurement requires high accuracy, therefore it is mandatory to control all the systematic uncertainties related to the reactor $\bar{\nu}_{e}$ source. For this purpose, the Taishan Antineutrino Observatory (TAO) TAO detector will measure with unprecedented energy resolution the non-oscillated $\bar{\nu}_{e}$ spectrum at $\sim 30$ m from one of the reactors exploited by JUNO. However, since the reactor $\bar{\nu}_{e}$ spectrum changes as a function of reactor fuel burnup, computational tools are needed to assess the $\bar{\nu}_{e}$ flux emitted by all reactors.
Until now, two different methods have been used to model the reactor $\bar{\nu}_{e}$ flux: the conversion Huber2011 ; Mueller2011 and the summation (also called ab-initioFallot2012 ; Estienne_ab_initio methods. Both approaches are affected by uncertainties related to the input data, leading to a discrepancy in both $\bar{\nu}_{e}$ flux intensity and spectral shape HayesVogel2016 ; zhang2023reactor when compared with the experimental data acquired by near detectors DayaBay:2022jnn ; DoubleChooz:2019qbj ; RENO:2015ksa ; Stereo2023 ; Ko_2017 ; Andriamirado_2023 .
@@
    the concentration of $\beta^{-}$ decaying nuclides in the reactor core as a function of time;
@@
In this paper we present the development of an analysis tool designed explicitly for probing the uncertainties inherent in the first ingredient. We choose as test model a standard PWR system due to the data availability of most of its components. Even if the reactor design is not strictly equal to the reactors placed in Taishan and Yangjiang, the presented results are representative of a light water reactor burnup cycle. This work is meant to show the methodological approach to reproduce the antineutrino spectrum from a reactor model, highlighting some features that can have an impact on it. Future works will be devoted to have an high fidelity simulation of both the plants involved in JUNO and TAO. Through meticulous simulation of reactor neutronics and fuel burnup, we attain a comprehensive understanding of the production and decay processes of all contributing nuclides that shape the $\bar{\nu}_{e}$ flux. In a second step we couple reactor burnup simulations to an extensive $\bar{\nu}_{e}$ spectral database to investigate the reactor-related uncertainties affecting the non-oscillated reactor $\bar{\nu}_{e}$ spectra. This aspect can be considered of extreme importance for all reactor neutrino experiments since it gives the key to reactor $\bar{\nu}_{e}$ flux control, allowing its real time spectral reconstruction and providing a continuous monitor of the $\bar{\nu}_{e}$ flux variations.
@@
  • in Sect. 4 we compare the fission fractions and the non oscillated reactor $\bar{\nu}_{e}$ spectrum at different burnup levels obtained with different approximations in the reactor simulations.
@@ 2 Methodology and tools
[…]
@@
$S(E_{\nu},t)=\frac{W_{\text{th}}(t)}{\sum_{i=1}^{4}E_{i}f_{i}(t)}\sum_{i=1}^{4}f_{i}(t)S_{i}(E_{\nu},t)$ (2)
where $i$ labels the main fissile nuclides in PWR reactors (${}^{235}$, ${}^{238}$, ${}^{239}$u, ${}^{241}$u), $W_{th}(t)$ is the reactor thermal power, $E_{i}$ is the average energy deposited in the reactor per fission of the $i$-th nuclide, $f_{i}(t)$ are the fission fractions (i.e. the ratio of the fission rate of the $i$-th nuclide to the total fission rate). Finally, $S_{i}(E_{\nu},t)$ is the spectrum associated to the fission of the $i$-th nuclide estimated with the summation method, where the antineutrino spectra of the daughters of $i$ are summed together, each with a weight that reflects their decay rate in the system. It is worth to highlight that the time dependence in $S_{i}(E_{\nu},t)$ will not be considered in this work because of the equilibrium hypothesis (see Sec. 2.2)
Equation (2) highlights that the reactor $\bar{\nu}_{e}$ spectrum depends on two sets of parameters: the first one, including $f_{i}(t)$ and $W_{\text{th}}(t)$, is related to reactor history, power and fuel burnup. The second comprises the nuclear data used to extract the different $\bar{\nu}_{e}$ spectra from the corresponding $\beta^{-}$ decays. In the following subsections, we describe the tools that we use to
  • calculate reactor-related parameters (2.1) using the flexible Serpent environment for fissile system simulations;
  • build the $\bar{\nu}_{e}$ spectra of the four actinides (2.2) adopting the summation method, with the recently developed BetaShape code.
@@
Reactor spectrum S(E_ν,t) =Wthi=14Eifi(t) ∑_i=1^4 f_i (t) S_i(E_ν) Summation method S_i(E_ν) = ∑_j=1^N_i y_i,j   S_j(E_ν) Fission yields y_i,j Daughters $\bf{\bar{\nu}_{e}}$ spectrumS_j(E) BetaShape Nuclear Database Fission Fractions f_i(t) MC simulation: Serpent Input: Geometry/materials and power history $W_{th}$ Antineutrino dataReactor data
@@ 2.1 The Serpent tool for reactor simulations
    $\frac{\text{d}N(\mathbf{r},t)}{\text{d}t}=\left[\int_{0}^{\infty}\phi(\mathbf{r},E,t)\mathbb{X}(E,T)\,\text{d}E+\mathbb{D}\right]N(\mathbf{r},t)$ (4)
@@
$R=\frac{1}{V}\int_{V}\int_{4\pi}\int_{E_{g-1}}^{E_{g}}h(\mathbf{r},E)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}\mathbf{\Omega}\,\text{d}E$ (5)
@@
$f_{i}=\frac{\int_{V}\int_{0}^{\infty}n_{i}(\mathbf{r})\sigma_{f,i}(E,T)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}E}{\sum_{k=1}^{4}\int_{V}\int_{0}^{\infty}n_{k}(\mathbf{r})\sigma_{f,k}(E,T)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}E}$ (6)
@@ 2.2 Antineutrino spectra production
$S_{i}(E_{\nu},t)=\frac{\sum_{j=1}^{N_{i}}A_{i,j}(t)\,S_{j}(E_{\nu})}{\sum_{j=1}^{N_{i}}A_{i,j}(t)\int_{0}^{\infty}\,S_{j}(E_{\nu})\text{d}E_{\nu}}$ (7)
@@
$A_{i,j}\propto y_{i,j}$ (8)
where $y_{i,j}$ is the cumulative fission yield (CFY), i.e. the probability of production of nuclide $j$ per each fission of $i$. Moreover, it is worth noting that most part of $\beta^{-}$ decays with Q-values higher than 1.8 MeV are characterized by half-lives much shorter than the typical reactor steady-state operation time, thus the equilibrium condition is quickly reached for most nuclides emitting $\bar{\nu}_{e}$ detectable through the IBD reaction. The adoption of the equilibrium approximation greatly simplifies the computation of $S_{i}(E_{\nu})$, since the values of $y_{i,j}$ and their uncertainties are tabulated in nuclear databases for fission induced by thermal, fast and 14 MeV neutrons.
@@ 3 Validation of reactor simulations
[…]
@@
In Fig. 2 we show the position of the rods SF95 and SF97 in their assemblies and in Table 1 we summarize the available information about the position and the local burnup (estimated through the ${}^{148}$d method Nd148_method1 ; Nd148_method2 ) of the samples taken from the rods labelled with a progressive number from highest to lowest position.
Assembly Sample ID Axial position (cm) Burnup (MWd/kg)
$\bf{{}^{235}}$ (g/tonU)
$\bf{{}^{238}}$ (g/tonU)
$\bf{{}^{239}}$u (g/tonU)
$\bf{{}^{241}}$u (g/tonU)
NT3G23
SF95-1
4.12
14.30
2.674×$\times 10^{4}$
9.499×$\times 10^{5}$
4.227×$\times 10^{3}$
3.69×$\times 10^{2}$
SF95-2
20.2
24.35
1.927×$\times 10^{4}$
9.424×$\times 10^{5}$
5.655×$\times 10^{3}$
9.578×$\times 10^{2}$
SF95-3
72.2
35.53
1.326×$\times 10^{4}$
9.338×$\times 10^{5}$
6.194×$\times 10^{3}$
1.486×$\times 10^{3}$
SF95-4
200.2
36.69
1.23×$\times 10^{4}$
9.335×$\times 10^{5}$
6.005×$\times 10^{3}$
1.466×$\times 10^{3}$
SF95-5
340.2
30.40
1.544×$\times 10^{4}$
9.388×$\times 10^{5}$
5.635×$\times 10^{3}$
1.153×$\times 10^{3}$
NT3G24
SF97-1
0.40
17.69
2.347×$\times 10^{4}$
9.493×$\times 10^{5}$
3.844×$\times 10^{3}$
4.237×$\times 10^{2}$
SF97-2
19.1
30.73
1.571×$\times 10^{4}$
9.377×$\times 10^{5}$
5.928×$\times 10^{3}$
1.235×$\times 10^{3}$
SF97-3
46.8
42.16
1.030×$\times 10^{4}$
9.282×$\times 10^{5}$
6.217×$\times 10^{3}$
1.689×$\times 10^{3}$
SF97-4
168
47.03
8.179×$\times 10^{3}$
9.246×$\times 10^{5}$
6.037×$\times 10^{3}$
1.77×$\times 10^{3}$
SF97-5
276.7
47.25
7.932×$\times 10^{3}$
9.247×$\times 10^{5}$
5.976×$\times 10^{3}$
1.754×$\times 10^{3}$
SF97-6
339.7
40.79
1.016×$\times 10^{4}$
9.310×$\times 10^{5}$
5.677×$\times 10^{3}$
1.494×$\times 10^{3}$
Table 1: Data of axial positions and local burnup of the samples drilled from the SF95 and SF97 pins. The axial locations have to be intended as the distance from the top of the fuel, as reported in the original documentation. The local burnup was evaluated with the ${}^{148}\text{Nd}$ technique, which brings a ±$\pm 3\%$ uncertainty. Finally we report the measured concentrations for ${}^{235}$, ${}^{238}$, ${}^{239}$u and ${}^{241}$u in units of grams per uranium tons in the fresh fuel. Data from Takahama3 .

### v3

@@ Abstract
Nuclear reactors are a source of electron antineutrinos due to the presence of unstable fission products that undergo $\beta^{-}$ decay. They will be exploited by the JUNO experiment to determine the neutrino mass ordering and to get very precise measurements of the neutrino oscillation parameters. This requires the reactor antineutrino spectrum to be characterized as precisely as possible both through high resolution measurements, as foreseen by the TAO experiment, and detailed simulation models. In this paper we present a benchmark analysis utilizing Serpent Monte Carlo simulations in comparison with real pressurized water reactor spent fuel data. Our objective is to study the accuracy of fission fraction predictions as a function of different reactor simulation approximations. Then, using the BetaShape software, we construct reactor antineutrino spectrum using the summation method, thereby assessing the influence of simulation uncertainties on it.
@@ 1 Introduction
which has a $\bar{\nu}_{e}$ energy threshold of 1.8 MeV.
One of the main goals of the JUNO experiment is to determine the neutrino mass ordering by analyzing the pattern of $\bar{\nu}_{e}$ oscillations at medium baseline. This experimental measurement requires high accuracy, therefore it is mandatory to control all the systematic uncertainties related to the reactor $\bar{\nu}_{e}$ source. For this purpose, the Taishan Antineutrino Observatory (TAO) TAO detector will measure with unprecedented energy resolution the non-oscillated $\bar{\nu}_{e}$ spectrum at $\sim 40$ m from one of the reactors exploited by JUNO. The importance of computational tools for predicting reactor antineutrino fluxes stands out in this framework, since in principle, if they are validated with experimental data, they can be used for flux shapes predictions, by knowing only reactor operation information. That could be the case for both the second Taishan reactor core as well as the six reactors of Yangjiang plant, which lack of near field detectors. Also, nuclear database accuracy could be benchmarked against high energy resolution TAO data.
Until now, two different methods have been used to model the reactor $\bar{\nu}_{e}$ flux: the conversion Huber2011 ; Mueller2011 and the summation (also called ab-initioFallot2012 ; Estienne_ab_initio methods. Both approaches are affected by uncertainties related to the input data, leading to a discrepancy in both $\bar{\nu}_{e}$ flux intensity and spectral shape HayesVogel2016 ; ZHANG2024104106 when compared with the experimental data acquired by near detectors DayaBay:2022jnn ; DoubleChooz:2019qbj ; RENO:2015ksa ; Stereo2023 ; Ko_2017 ; Andriamirado_2023 .
@@
    the activity of $\beta^{-}$ decaying nuclides in the reactor core as a function of time;
@@
In this paper we present the development of an analysis tool designed explicitly for probing the uncertainties inherent in the first ingredient. We choose as test model a standard PWR system due to the data availability of most of its components. Even if the reactor design is not strictly equal to the reactors placed in Taishan and Yangjiang, the presented results are representative of a light water reactor burnup cycle. In particular, the main differences in the present work are a different 235U assembly enrichment and a lower nominal power with respect to the two Chinese plants. This work is meant to show the methodological approach to reproduce the antineutrino spectrum from a reactor model, highlighting some features that can have an impact on it. Future works will be devoted to have a high fidelity simulation of both the plants involved in JUNO-TAO. Through meticulous simulation of reactor neutronics and fuel burnup, we attain a comprehensive understanding of the production and decay processes of all contributing nuclides that shape the $\bar{\nu}_{e}$ flux. In a second step we couple reactor burnup simulations to an extensive $\bar{\nu}_{e}$ spectral database to investigate the reactor-related uncertainties affecting the non-oscillated reactor $\bar{\nu}_{e}$ spectra. This aspect can be considered of extreme importance for all reactor neutrino experiments since it gives the key to reactor $\bar{\nu}_{e}$ flux control, allowing its real time spectral reconstruction and providing a continuous monitoring of the $\bar{\nu}_{e}$ flux variations.
@@
  • in Sect. 4 we compare the fission fractions and the non oscillated reactor $\bar{\nu}_{e}$ spectrum at different burnup levels obtained with different approximations in the reactor simulations;
  • in Sect. 5, summary and future works are pointed out.
@@ 2 Methodology and tools
During the operational phase of a Pressurized Water Reactor (PWR), thermal power $W_{th}$ is continuously produced through fission reactions. In the fresh fuel, usually made of uranium dioxide (UO2) enriched in 235U between 2 wtand 5 wt%, only uranium isotopes (235U and 238U) can be fissioned by neutrons. However, as reactor operation proceeds in time, a series of neutron capture reactions and radioactive decays leads to accumulation of 239Pu and 241Pu, which are fissile by thermal neutrons. The fission fragments are nuclei characterized by an excess of neutrons, thus they undergo a series of $\beta^{-}$ decays to reach nuclear stability. On average, about 6 $\bar{\nu}_{e}$ are emitted by $\beta^{-}$ decays after each fission event. It is therefore possible to consider each fissile element in the reactor as the parent of several unstable fission products.
@@
$S(E_{\nu},t)=\frac{W_{\text{th}}(t)}{\sum_{i=1}^{4}E_{i}f_{i}(t)}\sum_{i=1}^{4}f_{i}(t)S_{i}(E_{\nu},t),$ (2)
where $i$ labels the main fissile nuclides in PWR reactors (235U, 238U, 239Pu, 241Pu), $W_{th}(t)$ is the reactor thermal power, $E_{i}$ is the average energy deposited in the reactor per fission of the $i$-th nuclide, $f_{i}(t)$ are the fission fractions (i.e. the ratio of the fission rate of the $i$-th nuclide to the total fission rate). Finally, $S_{i}(E_{\nu},t)$ is the spectrum associated to the fission of the $i$-th nuclide estimated with the summation method, where the antineutrino spectra of the daughters of $i$ are summed together, each with a weight that reflects their decay rate in the system. It is worth to highlight that the time dependence in $S_{i}(E_{\nu},t)$ will not be considered in this work because of the equilibrium hypothesis (see Sec. 2.2). This choice is made to underline the effects of reactor parameters on the fission fractions, according to the most used approach in the international community adopting ab-initio scheme. Non-equilibrium studies will be faced in future works.
Equation (2) highlights that the reactor $\bar{\nu}_{e}$ spectrum depends on two sets of parameters: the first one, including $f_{i}(t)$ and $W_{\text{th}}(t)$, is related to reactor history, power and fuel burnup. The second comprises the nuclear data used to extract the different $\bar{\nu}_{e}$ spectra from the corresponding $\beta^{-}$ decays. In the following subsections, we describe the tools that we use to:
  • calculate reactor-related parameters (2.1) using the flexible Serpent (version 2.1.31) environment for fissile system simulations;
  • build the $\bar{\nu}_{e}$ spectra of the four actinides (2.2) adopting the summation method, with the recently developed BetaShape (version 2.3) code.
@@ 2.1 The Serpent tool for reactor simulations
    $\frac{\text{d}N(\mathbf{r},t)}{\text{d}t}=\left[\int_{0}^{\infty}\phi(\mathbf{r},E,t)\mathbb{X}(E,T)\,\text{d}E+\mathbb{D}\right]N(\mathbf{r},t),$ (4)
@@
$R=\frac{1}{V}\int_{V}\int_{4\pi}\int_{E_{g-1}}^{E_{g}}h(\mathbf{r},E)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}\mathbf{\Omega}\,\text{d}E,$ (5)
@@
$f_{i}=\frac{\int_{V}\int_{0}^{\infty}n_{i}(\mathbf{r})\sigma_{f,i}(E,T)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}E}{\sum_{k=1}^{4}\int_{V}\int_{0}^{\infty}n_{k}(\mathbf{r})\sigma_{f,k}(E,T)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}E},$ (6)
@@ 2.2 Antineutrino spectra production
$S_{i}(E_{\nu},t)=\frac{\sum_{j=1}^{N_{i}}A_{i,j}(t)\,S_{j}(E_{\nu})}{\sum_{j=1}^{N_{i}}A_{i,j}(t)\int_{0}^{\infty}\,S_{j}(E_{\nu})\text{d}E_{\nu}},$ (7)
@@
$A_{i,j}\propto y_{i,j},$ (8)
where $y_{i,j}$ is the cumulative fission yield (CFY), i.e. the probability of production of nuclide $j$ per each fission of $i$. Moreover, it is worth noting that most part of $\beta^{-}$ decays with Q-values higher than 1.8 MeV are characterized by half-lives much shorter than the typical reactor steady-state operation time, thus the equilibrium condition is quickly reached for most nuclides emitting $\bar{\nu}_{e}$ detectable through the IBD reaction. The adoption of the equilibrium approximation greatly simplifies the computation of $S_{i}(E_{\nu})$, since the values of $y_{i,j}$ and their uncertainties are tabulated in nuclear databases for fission induced by thermal, and fast neutrons. The correlations affecting the fission yield uncertainties are not included in this analysis since a complete database is currently missing.
@@ 3 Validation of reactor simulations
The Takahama-3 is a PWR operated by the Kansai Electric Power Company in Japan since 1984 IAEA_reactor . This reactor is characterized by a nominal thermal power of 2652 MW. Its core comprises 157 fuel assemblies, each of which contains a lattice structure of 17$\times$17 rod bundle. Each rod has a radius of $\sim 0.4$ cm, while the side of the squared lattice is 21.4 cm and the assembly total height is about 4 m with $\sim 3.64$ m of active length. Each fuel assembly contains 250 pins with standard fuel (i.e. UO2 with an enrichment of 235U/U = 4.11 wt%), 14 fuel pins with gadolinium as a burnable absorber (235U/U = 2.63 wtand a poison ratio of Gd2O3/Fuel = 6.0 wt%), and 25 guide tubes filled with water in correspondence to the control rod positions. In particular, spent nuclear fuel analyses were performed on three fuel rods, named SF95, SF96 and SF97, respectively belonging to the NT3G23 assembly (the first two) and to the NT3G24 assembly (the third one). SF95 and SF97 are standard fuel pins, whereas SF96 is a Gd-loaded fuel pin. These rods were irradiated in the core for a certain (documented) time; then, after the extraction, five or six samples were drilled out of them at different axial locations and analyzed with Isotope Diluted Mass Spectrometry (IDMS) to measure the concentration of uranium and plutonium isotopes. For this analysis, we chose to focus only on the four actinides that build up the fission fractions, since they are the most relevant isotopes in the reactor antineutrino framework. Moreover, since the Gd-loaded fuel exhibit a complex evolution due to a non-uniform depletion in the radial direction radialGd , and given that the available data refer exclusively to the axial direction, we decided to discuss only the data linked to SF95 and SF97 standard fuel pins, even if we aknowledge that it is an important point to be addressed for EPR and CPR reactors, which include Gd-fuel.
@@
In Fig. 2 we show the position of the rods SF95, SF96 and SF97 in their assemblies and in Table 1 we summarize the available information about the position and the local burnup (estimated through the 148Nd method Nd148_method1 ; Nd148_method2 ) of the samples taken from the rods labelled with a progressive number from highest to lowest position.
Assembly Sample ID Axial position (cm) Burnup (MWd/kg) 235U (g/tonU) 238U (g/tonU) 239Pu (g/tonU) 241Pu (g/tonU)
NT3G23 SF95-1 4.2 14.30 2.674×$\times 10^{4}$ 9.499×$\times 10^{5}$ 4.227×$\times 10^{3}$ 3.69×$\times 10^{2}$
SF95-2 20.2 24.35 1.927×$\times 10^{4}$ 9.424×$\times 10^{5}$ 5.655×$\times 10^{3}$ 9.578×$\times 10^{2}$
SF95-3 72.2 35.52 1.326×$\times 10^{4}$ 9.338×$\times 10^{5}$ 6.194×$\times 10^{3}$ 1.486×$\times 10^{3}$
SF95-4 200.2 36.69 1.23×$\times 10^{4}$ 9.335×$\times 10^{5}$ 6.005×$\times 10^{3}$ 1.466×$\times 10^{3}$
SF95-5 340.2 30.40 1.544×$\times 10^{4}$ 9.388×$\times 10^{5}$ 5.635×$\times 10^{3}$ 1.153×$\times 10^{3}$
NT3G24 SF97-1 0.40 17.69 2.347×$\times 10^{4}$ 9.493×$\times 10^{5}$ 3.844×$\times 10^{3}$ 4.237×$\times 10^{2}$
SF97-2 19.1 30.73 1.571×$\times 10^{4}$ 9.377×$\times 10^{5}$ 5.928×$\times 10^{3}$ 1.235×$\times 10^{3}$
SF97-3 46.8 42.16 1.030×$\times 10^{4}$ 9.282×$\times 10^{5}$ 6.217×$\times 10^{3}$ 1.689×$\times 10^{3}$
SF97-4 168 47.03 8.179×$\times 10^{3}$ 9.246×$\times 10^{5}$ 6.037×$\times 10^{3}$ 1.77×$\times 10^{3}$
SF97-5 276.7 47.25 7.932×$\times 10^{3}$ 9.247×$\times 10^{5}$ 5.976×$\times 10^{3}$ 1.754×$\times 10^{3}$
SF97-6 339.7 40.79 1.016×$\times 10^{4}$ 9.310×$\times 10^{5}$ 5.677×$\times 10^{3}$ 1.494×$\times 10^{3}$
Table 1: Data of axial positions and local burnup of the samples drilled from the SF95 and SF97 pins. The axial locations have to be intended as the distance from the top of the fuel, as reported in the original documentation. The local burnup was evaluated with the ${}^{148}\text{Nd}$ technique, which brings a ±$\pm 3\%$ uncertainty. Finally we report the measured concentrations for 235U, 238U, 239Pu and 241Pu in units of grams per uranium tons in the fresh fuel. Data from Takahama3 .